OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!! If ∠c ≅ ∠d, m∠c = 5x - 7, and m∠d = 3x + 25, how many degrees are in ∠c?

OpenStudy (anonymous):

you can allways tell the answer by the first two angles

OpenStudy (binary3i):

C= D 5X-7=3X-25 2X=-18 X=-9 c=5x-7 c=5*(-9) -7 c=-52